# Alberta Wells Dataset: Pinpointing Oil and Gas Wells from Satellite Imagery

### v1

@@ 1 Introduction
Across the world, there are millions of abandoned oil and gas wells left to degrade by the companies or individuals that built them. No longer producing usable fossil fuels, these wells nonetheless have a significant impact on the environment, with many of them leaking significant quantities of methane, a powerful greenhouse gas, into the atmosphere. In Canada, an estimated 370,000 abandoned wells produce methane equivalent to half a million metric tons of CO2 annually [1, 2], while in the U.S. there are an estimated 4 million abandoned wells [1], releasing over five million metric tons of CO2 equivalent emissions per year. Abandoned wells also pose health and safety concerns, in particular by leaching toxic chemicals into the groundwater of surrounding communities [3].
It is possible to plug abandoned wells to mitigate the harms associated with them (with so-called “super-emitter” wells an especially high priority [4, 5]). However, a significant fraction of abandoned wells remain unknown. In Pennsylvania, as much as 90% of abandoned wells are estimated to be unrecorded [5]. In Canada, abandoned wells have been described as the most uncertain source of methane emissions nationally due to the poor quality of data surrounding them [1].
@@
In this paper, we present the first large-scale machine-learning dataset for pinpointing onshore wells oil and gas wells, encompassing abandoned, suspended, and active wells. Our main contributions are as follows:
  • We introduce the Alberta Wells Dataset, which includes information on over 200k abandoned, suspended, and active onshore wells with high-resolution satellite imagery.
@@
  • We evaluate a wide range of deep learning algorithms commonly used for similar tasks, finding promising performance but opportunities for significant improvement.
@@ 2 Previous Work
Table 1: Previous datasets in which remote sensing algorithms are applied to detect oil and gas wells. “N/A” is given for datasets which do not indicate the number of individual wells in the dataset.
@@
Alberta Well Dataset
213,447 94,343 3 Alberta, Canada
Planet
Labs
Hundreds of satellites continuously monitor the Earth’s surface, generating petabyte-scale remote sensing datasets [6]. With advancements in hardware, the quality of remote sensing images has significantly improved in terms of spatial and temporal resolution. High-quality remote sensing data are available through state-funded projects like Sentinel and Landsat, and more recently through private enterprises such as Planet Labs [16]. Increasingly, machine learning has been used to parse such raw data, including in a wide range of applications for tackling climate change [7]. Benchmark datasets in this area have included tasks in land use and land cover (LULC) estimation [17], crop classification [18, 19], species distribution modeling [20], and forest monitoring [21].
@@ 3 Alberta Wells Dataset
In this paper, we introduce the Alberta Wells Dataset for oil and gas well detection. The dataset is drawn from the province of Alberta, Canada, a region with the third-largest oil reserves in the world and a substantial number of oil and gas wells, many of which have been present for over a century. The entire province of Alberta (an area larger than the UK and Germany combined) encompasses a diverse range of geographical zones and is highly diverse for a landlocked region, including prairies, lakes, forests, and mountains. The dataset contains over 94,000 patches of satellite imagery acquired from Planet Labs [16], covering more than 213,000 individual wells. Each patch is annotated with labels for both segmentation and bounding box localization. The annotations are based on data from the Alberta Energy Regulator, quality-controlled by domain experts.
(a) Training set
(b) Validation set
(c) Test set
Figure 1: Distribution of the number of individual wells in positive samples from the dataset. We also include an equal number of images with no wells at all.
Table 2: Statistics of instances and wells represented across the Alberta Wells Dataset.
Split Count Patches Total Count Wells Patches Count Non-Wells Patches Count of Well Type in Wells Patches of Split
Abandoned Suspended Active
Train 167436 83718 83718 46342 47595 100294
Validation 9463 4731 4731 3166 2671 2406
Test 11789 5894 5894 4024 3609 3340
@@
To simulate real-world conditions, we ensure a varied density of wells per image, as highlighted in Figure 1. We also include satellite imagery patches with no wells present from areas nearby to areas with wells, ensuring no overlap between the samples. This balanced dataset maintains an equal distribution of well and non-well images. Table 2 details the total sample count in each dataset split, alongside the number of well and non-well patches.
3.1 Well Data Collection, Quality Control & Patch Creation
@@
The Alberta Energy Regulator (AER) oversees the energy industry in the province, ensuring companies adhere to regulations as they develop oil and gas resources. AER publishes AER ST37 [30], a monthly list of all wells reported in Alberta, detailing their geographic location, mode of operation, license status, and type of product being extracted, among other attributes. This data provides a metadata (.txt) file and a .shp shape file, where each entry represents a unique geo-location point per site but often contains duplicates. However, this data cannot be used directly because the license status or mode of operation does not always correlate with the actual status of the well and often contains duplicates. Therefore, we work with domain experts to perform quality control on the dataset as illustrated in Figure 2.
First, we remove duplicate entries from the well metadata, which often contain multiple instances of the same well identified by duplicate license numbers. We resolve these duplicates by retaining the most recent update. A similar approach is applied to the shapefile, where duplicates are resolved using the license date. Afterward, we merge both datasets and filter the data, categorizing the wells as active, abandoned, or suspended based on specific criteria developed in consultation with domain experts, as shown in Table 3. We check for duplicate location coordinates in the dataset and resolve them by retaining the instance with the latest drill date. Finally, we ensure all the well instances in the dataset are indeed within the boundaries of Alberta. The raw metadata file has around 637,000 instances, while the surface hole geometry file has around 512,000 instances. After quality control and filtering, we have around 217,000 instances.
After filtering and performing quality control on the datasets with domain experts, we calculate the geographical bounds covered by the well instances across the province and divide the region into non overlapping square image patches, each covering an area of 1.1025 sq km (with sides of 1050m). These images include various numbers of individual wells (see Fig. 1), and we ensure that an approximately equal number of patches exist with and without wells. As a result of this process, some samples were excluded due to being located outside Alberta’s geographical boundaries, leading to a final total of approximately 213,000 well instances in the dataset patches.
Figure 2: AER ST37 Dataset Cleaning and Quality Control
@@ 3.2 Dataset Splitting
To create a well-distributed dataset that represents various geographical regions and offers a diverse dataset for evaluation and testing, we developed a splitting algorithm (see Algorithm 1). Our splitting approach focuses on balancing regions, not individual examples, ensuring that both the training and test sets reflect a diverse range of regions from Alberta’s varied landscape. This approach preserves dataset diversity and simulates real-world conditions where imbalances are common.
(a) Step-1 k1i clusters
(b) c2i cluster centroids
(c) Step-2 k2i clusters
(d) Final Dataset Split
Figure 3: Illustration of the outcome of applying our dataset splitting algorithm: In Figures (a) to (c), different colors represent various cluster IDs. In Figure (d), blue refers to the training set, orange to the validation set, and green to the test set.
@@
This method involves forming small clusters $k_{1i}$ of nearby well patches based on their centroids as illustrated in Figure 3 (a). These small clusters are then grouped into larger, non-intersecting super-clusters $k_{2i}$, with each super-cluster representing a city or larger geographical area. The formation of super-clusters involves calculating a centroid for each $k_{1i}$ cluster based on the centroids of the well patches it contains as illustrated in Figure 3 (b). By clustering wells in this manner, we ensure that $k_{1i}$ clusters group wells from nearby localities together, while $k_{2i}$ clusters group wells from the same geographic region as illustrated in Figure 3 (c). Thus, each $k_{2i}$ cluster represents a geographic distribution, with each $k_{1i}$ cluster within it representing a sample of that distribution.
To ensure a diverse and well-distributed evaluation and testing of our machine learning model, we select the $k_{1i}$ clusters with the two fewest well instances from each $k_{2i}$ super-cluster for inclusion in the evaluation and test sets. This approach ensures a diverse representation of the dataset as observed in Figure 3 (d). Moreover, we maintain an equal distribution of well and non-well patches. In cases of imbalance in non-well images, we exclude such patches from the contributing $k_{1i}$ clusters as specified in Algorithm 1. For imbalances in well images, we sample non-well patches that are not part of any other clusters.
The parameters used in constructing the dataset are $M=300$ and $N=30$. These were picked heuristically so as to create a well-distributed dataset. Alberta’s varied landscape offers a rich environment for creating a comprehensive oil well dataset. Training machine learning models on this extensive dataset improves their robustness and ability to generalize to similar, less-studied regions, thereby supporting well detection and efforts to mitigate global warming. By forming non-overlapping clusters ($k_{1i}$), each with its own well and non-well patches, we minimize the risk of data leakage while ensuring diversity. We also balanced non-well images across clusters to better simulate real-world conditions. This approach helps maintain the diversity of the dataset.
@@ 3.3 Satellite Imagery Acquisition & Label Creation
Figure 4: A sample image patch from our dataset includes examples with no wells, two wells, and multiple wells. Additionally, we present qualitative results with predictions generated by our Segmentation U-Net (EfficientNet-B6) and Object Detection FCOS models.
@@
We used PlanetScope-4-Band imagery [16] featuring RGB and Near Infrared bands to represent satellite images of the region with a medium resolution of about 3 meters per pixel. PlanetScope, a product of Planet Labs, consists of approximately 130 satellites that can image the entire Earth’s land surface daily, collecting up to 200 million sq. km of data each day. We obtained Surface Reflectance imagery, which is offset-corrected, flat-field-corrected, ortho-rectified, visually processed, and radio-metrically corrected. These processes ensure consistency across varying atmospheric conditions and minimize uncertainty in spectral response over time and location, making the data ideal for temporal analysis and monitoring applications.
We choose Planet Labs data over other alternatives since it is updated daily, making it possible to pick a consistent time for all the images, which is important for training dataset consistency. It also provides multispectral imagery (4-band: RGB+Near Infrared), and the Near Infrared band is a useful addition since certain features, like ground depressions indicating well sites, may be more detectable in this band. Lastly, while other alternatives may be limited in remote regions, Planet’s global satellite constellation ensures more consistent coverage. All the imagery we use is made publicly available in the dataset.
To ensure the highest quality, we selected images with no cloud cover. The images were acquired by Planet satellites within a timeframe that aligns with the well-location data from AER. We obtained satellite images for each sample based on geographical coordinates, ensuring an intersection between the actual area of interest and the acquired imagery.
@@
We frame the task of identifying wells as both an object detection and segmentation task since related remote sensing tasks have found both framings to be constructive. For each image patch, as shown in Figure 4, we generated corresponding segmentation maps and object detection annotations for all known wells in the image based on the point labels provided in the AER data. For binary segmentation, we annotated each well site with a circle to match the teardrop shape typical for well sites. We standardized the diameter of a well site to a value of 90 meters (such sites typically range from 70 to 120 meters in diameter). We used the same scale to define bounding boxes in the object detection task, following the COCO [31] format for annotations. Additionally, we created multi-class segmentation maps, where each class represents a different state of the well (active, suspended, or abandoned), and included this information in the object detection annotations. (We do not perform multi-class segmentation experiments here, but it is possible that future researchers may find this task useful.)
4 Benchmark Experiments
@@
We train benchmark deep learning models for both the binary segmentation and object detection tasks. Our focus includes all oil and gas wells, regardless of their operational status, since they exhibit similar footprints and consistent features, making them detectable in satellite imagery.
@@
For both tasks, we augment images by randomly resizing images to 256$\times$256, ensuring all bounding boxes remain intact for object detection. We then apply horizontal and vertical flipping with a probability of 0.25 each, followed by normalization using channel-wise mean and standard deviation calculated from the training split of the dataset. The hyperparameters we use in these various models represent standard performant settings and are not intended to represent the outcome of hyperparameter optimization.
@@ 4.1 Binary Segmentation
We selected well-known baseline models for binary segmentation, encompassing the deep CNN-based approaches U-Net [32], PAN[33], and DeepLabV3+ [34] as well as the Transformer-based architectures Segformer [35] and UperNet [36].
U-Net [32] was chosen for its widespread use as a baseline, offering an effective encoder-decoder architecture for multi-scale feature extraction. PAN [33] improves multi-scale context with pyramid pooling and attention mechanisms. DeepLabV3+[34] was selected for its popularity in remote sensing tasks with its Atrous Convolution and ASPP module for capturing contextual information at various scales. SegFormer [35] is a transformer-based architecture designed for semantic segmentation, utilizing self-attention mechanisms for capturing long-range dependencies. UperNet [36] combines UNet [32] and PSPNet [37] architectures, featuring a UNet-like structure for multi-scale feature fusion and PSPNet’s pyramid pooling module integrated with a Swin Transformer [38] backbone for efficient multi-scale processing.
We train all CNN-based models using a ResNet50 [39] backbone, a batch size of 128, and the BCELogits loss function. To fine-tune the model, a cosine annealing scheduler [40] is used, which adjusts the learning rate smoothly in a cyclical manner by gradually decreasing it. To evaluate the impact of backbones with larger receptive fields and attention mechanisms, we also experimented with additional backbones with U-Net. This included ResNeXt50 [41], which enhances feature learning through grouped convolutions; SE-ResNet50 [42], which introduces channel-wise attention with Squeeze-and-Excitation blocks; and EfficientNetB6 [43], known for its balanced scaling. For transformer-based models, while both Segformer and UperNet use a Dice loss function and a polynomial learning rate scheduler, Segformer utilizes a mit-b0-ade [35] backbone with a batch size of 128, and UperNet employs a Swin Small Transformer with a batch size of 64. All models are optimized using AdamW [44] for 50 epochs.

### v2

@@ 1 Introduction
Across the world, there are millions of abandoned oil and gas wells left to degrade by the companies or individuals that built them. No longer producing usable fossil fuels, these wells nonetheless have a significant impact on the environment, with many of them leaking significant quantities of methane, a powerful greenhouse gas, into the atmosphere. In aggregate, these emissions represent the equivalent of millions of tons of carbon dioxide per year [1]. In Canada, an estimated 370,000 abandoned wells produce methane equivalent to half a million metric tons of CO2 annually [1, 2], while in the U.S., there are an estimated 4 million abandoned wells [1], releasing over five million metric tons of CO2 equivalent emissions per year.
Despite the small percentage, the high global warming potential of methane underscores the need for improved monitoring and management to reduce environmental impact. Abandoned wells also pose health and safety concerns, in particular by leaching toxic chemicals into the groundwater of surrounding communities [3]. It is possible to plug abandoned wells to mitigate the harms associated with them (with so-called “super-emitter” wells an especially high priority [4, 5]). However, a significant fraction of abandoned wells remain unknown. In Pennsylvania, as much as 90% of abandoned wells are estimated to be unrecorded [5]. In Canada, abandoned wells have been described as the most uncertain source of methane emissions nationally due to the poor quality of data surrounding them [1].
@@
In this work, we present the first large-scale machine-learning benchmark dataset for pinpointing onshore oil and gas wells, encompassing abandoned, suspended, and active wells. Our main contributions are as follows:
  • We introduce the Alberta Wells Dataset, containing information on over 200k abandoned, suspended, and active onshore oil and gas wells with high-resolution satellite imagery.
@@
  • We evaluate a wide range of state-of-the-art deep learning algorithms, showing promising performance but emphasizing the challenging nature of the task.
@@ 2 Previous Work
Hundreds of satellites continuously monitor the Earth’s surface, generating petabyte-scale remote sensing datasets [6]. With advancements in hardware, the quality of remote sensing images has significantly improved in terms of spatial and temporal resolution. High-quality remote sensing data are available through state-funded projects like Sentinel and Landsat, and more recently through private enterprises such as Planet Labs [9].
Increasingly, machine learning has been used to parse such raw data, including in a wide range of applications for tackling climate change [7]. Benchmark datasets in this area have included tasks in land use and land cover (LULC) estimation [10], crop classification [11, 12], species distribution modeling [13], and forest monitoring [14].Some datasets like SpaceNet 7 [15] include a few cases of study sites with oil wells, although the dataset was developed for multi-temporal urban monitoring.
Table 1: Previous datasets in which remote sensing algorithms are applied to detect oil and gas wells. “N/A” is given for datasets that do not indicate the number of individual wells in the dataset.
@@
Alberta Well Dataset (Ours) 213,447 94,343 3 Alberta, Canada Planet Labs
Table 2: Statistics of instances and wells across the dataset.
Split Total Patches Wells/ Non-Wells Patches Count of Well Type in Wells Patches of Split
Abandoned Suspended Active
Train 167436 83718 46342 47595 100294
Validation 9463 4731 3166 2671 2406
Test 11789 5894 4024 3609 3340
@@ 3 Alberta Wells Dataset
We introduce the Alberta Wells Dataset for oil and gas well detection. The dataset is drawn from the province of Alberta, Canada, a region with the third-largest oil reserves in the world and a substantial number of oil and gas wells, many of which have been present for over a century. The entire province of Alberta (an area larger than the UK and Germany combined) encompasses a diverse range of geographical zones and is highly diverse for a landlocked region, including prairies, lakes, forests, and mountains. The dataset contains over 188,000 patches of satellite imagery acquired by Planet Labs [9], which we make publicly available, with around 94,000 containing wells covering more than 213,000 individual wells. Each patch is annotated with labels for both segmentation and bounding box localization. The annotations are based on data from the Alberta Energy Regulator [31], quality-controlled by domain experts.
@@
To simulate real-world conditions, we ensure a varied density of wells per image, as highlighted in Figure LABEL:fig:figure1. We also include satellite imagery patches with no wells present from areas nearby to areas with wells, ensuring no overlap between the samples. This balanced dataset maintains an equal distribution of well and non-well images. Table 2 details the total sample count in each dataset split. It also mentions the count of oil wells in active, abandoned, and suspended states in each dataset split, emphasizing the dataset’s comprehensive nature.
3.1 Data Collection, Quality Control & Patch Creation
The Alberta Energy Regulator (AER) oversees the energy industry in the province, ensuring companies adhere to regulations as they develop oil and gas resources. AER publishes AER ST37 [31], a monthly list of all wells reported in Alberta, detailing their geographic location, mode of operation, license status, and type of product being extracted, among other attributes. This data provides a metadata (.txt) file and a .shp shape file, where each entry represents a unique geo-location point per site but often contains duplicates. However, this data cannot be used directly because the license status or mode of operation does not always correlate with the actual status of the well and often contains duplicates. Therefore, we work with domain experts to perform quality control as illustrated in Figure 3.
First, we remove duplicate entries from the well metadata, which often contain multiple instances of the same well identified by duplicate license numbers. We resolve these duplicates by retaining the most recent update. A similar approach is applied to the shapefile, where duplicates are resolved using the license date.
Afterward, we merge both datasets and filter the data, categorizing the wells as active, abandoned, or suspended based on criteria developed in consultation with domain experts, as shown in Table 3. We check for duplicate location coordinates in the dataset and resolve them by retaining the instance with the latest drill date.
Finally, we ensure all the well instances in the dataset are within Alberta’s boundaries. The raw metadata file has around 637,000 instances, while the surface hole geometry file has around 512,000 instances. After filtering and performing quality control on the datasets with domain experts, we have around 217,000 instances. We then calculate the geographical bounds covered by the well instances across the province and divide the region into non-overlapping square image patches, each covering an area of 1.1025 sq km (with sides of 1050m).
These images include various numbers of individual wells (see Fig. LABEL:fig:figure1), and we ensure that an approximately equal number of patches exist with and without wells. As a result of this process, some samples were excluded due to being located outside Alberta’s geographical boundaries, leading to a final total of approximately 213,000 well instances in the dataset patches.
Figure 3: Flowchart depicting our process for AER ST37 dataset cleaning and quality control.
@@ 3.2 Dataset Splitting
To create a well-distributed dataset that represents various geographical regions and offers a diverse dataset for evaluation and testing, we developed a splitting algorithm (see Algorithm 1). Our splitting approach focuses on balancing regions, not individual examples, ensuring that both the training and test sets reflect a diverse range of regions from Alberta’s varied landscape. This approach preserves diversity and simulates real-world conditions where imbalances are common.
This method involves forming small clusters $k_{1i}$ of nearby well patches based on their centroids as illustrated in Figure LABEL:fig:figuresplit (a). These small clusters are then grouped into larger, non-intersecting super-clusters $k_{2i}$, with each super-cluster representing a city or larger geographical area. The formation of super-clusters involves calculating a centroid for each $k_{1i}$ cluster based on the centroids of the well patches it contains as illustrated in Figure LABEL:fig:figuresplit (b). By clustering wells in this manner, we ensure that $k_{1i}$ clusters group wells from nearby localities together, while $k_{2i}$ clusters group wells from the same geographic region as illustrated in Figure LABEL:fig:figuresplit (c). Thus, each $k_{2i}$ cluster represents a geographic distribution, with each $k_{1i}$ cluster within it representing a sample of that distribution.
To ensure a diverse and well-distributed evaluation and testing of our machine learning model, we select the $k_{1i}$ clusters with the two fewest well instances from each $k_{2i}$ super-cluster for inclusion in the evaluation and test sets. This approach ensures a diverse representation of the dataset as observed in Figure LABEL:fig:figuresplit (d).
@@
Moreover, we maintain an equal distribution of well and non-well patches. In cases of imbalance in non-well images, we exclude such patches from the contributing $k_{1i}$ clusters as specified in Algorithm 1. For imbalances in well images, we sample non-well patches that are not part of any other clusters.
The parameters used in constructing the dataset are $M=300$ and $N=30$. These were picked heuristically so as to create a well-distributed dataset. Alberta’s varied landscape offers a rich environment for creating a comprehensive oil well dataset. Training machine learning models on this extensive dataset improves their robustness and ability to generalize to similar, less-studied regions, thereby supporting well detection and efforts to mitigate global warming. By forming non-overlapping clusters ($k_{1i}$), each with its own well and non-well patches, we minimize the risk of data leakage while ensuring diversity. We also balanced non-well images across clusters to better simulate real-world conditions and maintain the diversity of the dataset.
@@ 3.3 Satellite Imagery Acquisition & Label Creation
We used PlanetScope-4-Band imagery [9] featuring RGB and Near Infrared bands to represent satellite images of the region with a medium resolution of about 3 meters per pixel. PlanetScope, a product of Planet Labs, consists of approximately 130 satellites that can image the entire Earth’s land surface daily, collecting up to 200 million sq. km of data each day. We obtained Surface Reflectance imagery, which is offset-corrected, flat-field-corrected, ortho-rectified, visually processed, and radio-metrically corrected. These processes ensure consistency across varying atmospheric conditions and minimize uncertainty in spectral response over time and location, making the data ideal for various applications. We choose Planet Labs data over other alternatives since it is updated daily, making it possible to pick a consistent time for all the images, which is important for training dataset consistency. It also provides multispectral imagery (4-band: RGB+Near Infrared), and the Near Infrared band is a useful addition (as we later verify in Table 6) since certain features, like ground depressions indicating well sites, may be more detectable in this band. Lastly, while other alternatives may be limited in remote regions, Planet’s global satellite constellation ensures more consistent coverage. All the imagery we use is made publicly available in the dataset. To ensure the highest quality, we selected images with no cloud cover. The images were acquired by Planet satellites within a timeframe that aligns with the well-location data from AER. We obtained satellite images for each sample based on geographical coordinates, ensuring an intersection between the actual area of interest and the acquired imagery.
We frame the task of identifying wells as both an object detection and segmentation task. In a real-world setting, high performance in either detection or segmentation would be sufficient for detection of wells; therefore we test both framings since it is not a priori clear which may be best suited to remote sensing algorithms. For each image patch, as shown in Figure 7, we generated corresponding segmentation maps and object detection annotations for all known wells in the image based on the point labels provided in the AER data.
@@
For binary segmentation, we annotated each well site with a circle to match the teardrop shape typical for well sites. We standardized the diameter of a well site to a value of 90 meters (such sites typically range from 70 to 120 meters in diameter). We used the same scale to define bounding boxes in the object detection task, following the COCO [32] format for annotations. The overlap in ground truth bounding boxes for some of the wells in Figure 7 and Figure 14 reflects the clustering of multiple wells in densely developed oil and gas sites, where the spatial overlap of wells and infrastructure is common. (Note that this is a characteristic of the data, not a limitation of our quality control strategy.) Additionally, we created multi-class segmentation maps, where each class represents a different state of the well (active, suspended, or abandoned), and included this information in the object detection annotations. (We do not perform multi-class segmentation experiments here, but it is possible that future researchers may find this task useful.)
Identifying wells at 3m resolution seems like a very difficult task That said, it’s important to consider that ML models may actually have the potential to surpass human performance in this particular task, for several reasons. First, identifying wells at a 3m resolution from satellite imagery is extremely challenging and requires significant patience by a human. Additionally, there are few human experts in this field, as well identification typically relies on existing databases or on-the-ground field studies rather than detailed satellite image analysis. Finally, certain features (such as depressions in the ground representing well sites) may be easier to pick out in the infrared channel, which humans are not accustomed to using, as compared with the RGB channels.
4 Experiments
@@
We evaluate the performance of deep learning models for binary segmentation and object detection tasks. Our focus includes all oil and gas wells, regardless of their operational status, since they exhibit similar footprints and consistent features, making them detectable in satellite imagery.
@@
We augment images by randomly resizing images to 256$\times$256, ensuring all bounding boxes remain intact for object detection. We then apply horizontal and vertical flipping with a probability of 0.25 each, followed by normalization using channel-wise mean and standard deviation calculated from the training split of the dataset.
The hyperparameters we use in these various models represent standard performant settings and are not intended to represent the outcome of hyperparameter optimization.
@@ 4.1 Binary Segmentation
We selected well-known baseline models for binary segmentation, encompassing the deep CNN-based approaches U-Net, PAN, and DeepLabV3+, as well as the Transformer-based architectures Segformer and UperNet.
U-Net [33] was chosen for its widespread use as a baseline, offering an effective encoder-decoder architecture for multi-scale feature extraction. PAN [34] improves multi-scale context with pyramid pooling and attention mechanisms. DeepLabV3+ [35] was selected for its popularity in remote sensing tasks with its Atrous Convolution and ASPP module for capturing contextual information at various scales. SegFormer [36] is a transformer-based architecture designed for semantic segmentation, utilizing self-attention mechanisms for capturing long-range dependencies. UperNet [37] combines UNet and PSPNet [38] architectures, featuring a UNet-like structure for multi-scale feature fusion and PSPNet’s pyramid pooling module integrated with a Swin Transformer [39] backbone for efficient multi-scale processing.
We train all CNN-based models using a ResNet50 backbone, a batch size of 128, and the BCELogits loss function. To fine-tune the model, a cosine annealing scheduler [40] is used, which adjusts the learning rate smoothly in a cyclical manner by gradually decreasing it. We also experimented with additional backbones with U-Net to evaluate the impact of backbones with larger receptive fields and attention mechanisms. This included ResNeXt50 [41], which enhances feature learning through grouped convolutions; SE-ResNet50 [42], which introduces channel-wise attention with Squeeze-and-Excitation blocks; and EfficientNetB6 [43], known for its balanced scaling.
In segmentation, the combination of UperNet’s contextual aggregation and the ConvNeXt backbone’s efficient representation enables precise segmentation, particularly valuable for remote sensing applications. For transformer-based models, while both Segformer and UperNet use a Dice loss function and a polynomial learning rate scheduler, Segformer utilizes a mit-b0-ade [36] backbone with a batch size of 128, while UperNet uses ConvNeXT (small and base) and Swin Transformer backbones with a batch size of 64. All models are optimized using AdamW for 50 epochs.